# A new expansion of the Riemann zeta function

### v3

@@ Abstract
We introduce a family of orthogonal polynomials $P_{m}(s)$, whose roots lie on the line $\{\Re(s)=1/2\}$, and we give an expansion of the Riemann zeta function in the critical strip as a convergent series $\sum_{m\geq 0}x_{m}P_{m}(s)$. The coefficients $x_{m}$of this expansion are expressed as linear combinations with rational coefficients of the Euler constant $\gamma$ and the values $\zeta(2),$ $\zeta(3),$ $\dots,$ $\zeta(m+1).$
@@ 1 Introduction
Some polynomial expansions of the Riemann zeta function (cf.[E-K-K], [F-V],[B],[M], [V]) are based on the Newton interpolation formula for functions related to Zeta. For example
@@
The polynomial expansion of the Riemann zeta function that we give in this paper is related to the Mellin transform of the functions $x\mapsto e^{-x}L_{m}(2x)$ where $(L_{m})_{m\geq 0}$ is the sequence of Laguerre polynomials defined by
@@
then, after an explicit evaluation of $z_{n}$, we obtain, for $0<\Re(s)<1$, the following expansion
$\zeta(s)=\sum_{m=0}^{+\infty}x_{m}P_{m}(s)$
@@ 2.1 Ramanujan’s constant of a series
Let’s assume that the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\ dx$ is convergent. The previous formula gives
@@
The constant $C(f)$ is called Ramanujan’s constant of the series $\sum_{n\geq 1}f(n)$ (cf. [H1]).
@@
and we see that the Ramanujan’s constant of the series $\sum_{n\geq 1}\frac{1}{n}$ is the Euler constant :
@@ 2.2 Plana’s formula
Thus, in this case the Ramanujan’s constant $C(f)$, previously defined in (2), is also
@@
Let $f$ and $g$ two functions verifying the assumptions of theorem 1 and such that $f(n)=g(n)$ for every integer $n\geq 1$. Then Ramanujan’s constants $C(f)$ and $C(g)$, defined by (4), need not necessarily be the same. For example, if $f\equiv 0$ and $g:z\mapsto\sin(\pi z)$, then $C(f)=0$, but
@@
To avoid functions like $z\mapsto\sin(\pi z)$, we will consider a space of analytic functions such that Ramanujan’s constant, defined by (4), depends only on the values $f(n)$ for all integers $n\geq 1$}.
@@ 2.3 The space $\mathcal{E}$
Let $\mathcal{E}$} the space of analytic functions $f$ in the half plane $\{\Re(z)>0\}$ such that there is some constants $\emph{a}<\pi$ and $\emph{b}>0$ with :
@@
$|g(x+it)|\sim\frac{1}{2}e^{\pi|t|}\text{ when }t\to\infty$
@@ 2.5 Interpolation of partial sums
To evaluate Ramanujan’s summation of a divergent series $\sum_{n\geq 1}f(n)$, a simpler way than the formula (6) can be obtained using an interpolation function of the partial sums of this series. In chapter VI of his Notebooks, Ramanujan uses, without any definition, such a function $\varphi$ which satisfies
@@
Let $f\in\mathcal{E}$. There is a unique function $\varphi_{f}\in\mathcal{E}$, such that
@@ 3.5 Orthogonality relations
$\Im\zeta(\frac{1}{2}+it)$ $=$ $2\sum_{m=0}^{+\infty}x_{2m+1}Q_{2m+1}(t).$

### v4

# arXiv:2512.11405v4[math.NT] 01 Mar 2026

@@ Abstract
We introduce a family of orthogonal polynomials $P_{m}(s)$, whose roots lie on the line $\{\Re(s)=1/2\}$, and we give an expansion of the Riemann’s zeta function in the critical strip as a convergent series $\sum_{m\geq 0}x_{m}P_{m}(s)$. The coefficients $x_{m}$of this expansion are expressed as linear combinations with rational coefficients of the Euler constant $\gamma$ and the values $\zeta(2),$ $\zeta(3),$ $\dots,$ $\zeta(m+1).$
@@ 1 Introduction
Some polynomial expansions of the Riemann’s zeta function (cf.[E-K-K], [F-V],[B],[M], [V]) are based on the Newton interpolation formula for functions related to the zeta function. For example
@@
The polynomial expansion of the Riemann’s zeta function that we give in this paper is related to the Mellin transform of the functions $x\mapsto e^{-x}L_{m}(2x)$ where $(L_{m})_{m\geq 0}$ is the sequence of Laguerre polynomials defined by
@@
then, after an explicit evaluation of $z_{m}$, we obtain, for $0<\Re(s)<1$, the following expansion
$\zeta(s)=\sum_{m=0}^{+\infty}x_{m}P_{m}(s),$
@@ 2.1 Ramanujan’s constant of a series
Let’s assume that the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\ dx$ is convergent. Then the previous formula gives
@@
The constant $C(f)$ is called the Ramanujan constant of the series $\sum_{n\geq 1}f(n)$ (cf. [H1]).
@@
and we see that the Ramanujan constant of the series $\sum_{n\geq 1}\frac{1}{n}$ is the Euler constant :
@@ 2.2 Plana’s formula
Thus, in this case the Ramanujan constant $C(f)$, previously defined in (2), is also
@@
Let $f$ and $g$ two functions verifying the assumptions of theorem 1 and such that $f(n)=g(n)$ for every integer $n\geq 1$. Then the Ramanujan constants $C(f)$ and $C(g)$, defined by (4), need not be the same. For example, if $f\equiv 0$ and $g:z\mapsto\sin(\pi z)$, then $C(f)=0$, but
@@
To avoid functions like $z\mapsto\sin(\pi z)$, we will consider a space of analytic functions such that the Ramanujan constant, defined by (4), only depends on the values $f(n)$ for all integers $n\geq 1$}.
@@ 2.3 The space $\mathcal{E}$
Let $\mathcal{E}$} the space of analytic functions $f$ in the half plane $\{\Re(z)>0\}$ such that there are some constants $\emph{a}<\pi$ and $\emph{b}>0$ with :
@@
$|g(x+it)|\sim\frac{1}{2}e^{\pi|t|}\text{ when }t\to\infty.$
@@ 2.5 Interpolation of partial sums
To evaluate Ramanujan’s summation of a divergent series $\sum_{n\geq 1}f(n)$, a simpler formula than (6) can be obtained by means of an interpolation function of the partial sums of this series. In chapter VI of his Notebooks, Ramanujan uses, without any definition, such a function $\varphi$ which satisfies
@@
Let $f\in\mathcal{E}$. There exists a unique function $\varphi_{f}\in\mathcal{E}$, such that
@@ 3.5 Orthogonality relations
$\Im\zeta(\frac{1}{2}+it)$ $=$ $-2i\sum_{m=0}^{+\infty}x_{2m+1}Q_{2m+1}(t).$